# BeyondMimic: From Motion Tracking to Versatile Humanoid Control via Guided Diffusion

### v1

@@ Abstract
Learning skills from human motions offers a promising path toward generalizable policies for whole-body humanoid control, yet two key cornerstones are missing: (1) a high-quality motion tracking framework that faithfully transforms large-scale kinematic references into robust and extremely dynamic motions on real hardware, and (2) a distillation approach that can effectively learn these motion primitives and compose them to solve downstream tasks. We address these gaps with BeyondMimic , the first real-world framework to learn from human motions for versatile and naturalistic humanoid control via guided diffusion. Our framework provides a motion tracking pipeline capable of challenging skills such as jumping spins, sprinting, and cartwheels with state-of-the-art motion quality. Moving beyond mimicking existing motions and synthesize novel ones, we further introduce a unified diffusion policy that enables zero-shot task-specific control at test time using simple cost functions. Deployed on hardware, BeyondMimic performs diverse tasks at test time, including waypoint navigation, joystick teleoperation, and obstacle avoidance, bridging sim-to-real motion tracking and flexible synthesis of human motion primitives for whole-body control. https://beyondmimic.github.io/.
@@ I Introduction
We introduce BeyondMimic, the first real-world framework to address the above two challenges altogether. BeyondMimic first introduces a motion tracking pipeline that is able to perform highly dynamic motions, such as jumping spins, sprinting, and cartwheels, on real humanoid hardware with state-of-the-art motion quality. Beyond simply mimicking these human motions on humanoids, our framework further presents a guided diffusion policy that synthesizes these motions or skills into a single unified policy that supports zero-shot, task-specific control at test time using only simple cost functions, solving downstream tasks flexibly without additional retraining.
@@
  • First Guided Diffusion for Humanoids: The first application of loss-guided diffusion to real-world humanoid whole-body control, enabling diverse downstream tasks via simple loss functions at test time.
@@ III Scaleable Motion Tracking
We propose a scalable pipeline for training motion-tracking policies. Using a single MDP configuration and a common set of hyperparameters, the pipeline produces high-quality sim-to-real policies given minutes-long references, demonstrating its scalability towards a diverse set of motions. We introduce the details below. Note that since our tracking framework does not have a history, we omit timestep $t$ in the formulation for clarity. Unless otherwise stated, all quantities in this section are expressed in the world frame.
@@ III-A Tracking Objective
The goal of our framework is to reproduce a given reference motion on robot hardware with high quality in global coordinates. However, the sim-to-real gap often leads to inevitable global drift. To preserve style while allowing global drift, the controller should not track absolute body poses. Instead, we designate a reference body $b_{\text{ref}}\in\mathcal{B}$, where $\mathcal{B}$ is the set of all robot bodies and $b_{\text{ref}}$ is typically the root or torso, as an anchor.
To anchor the desired pose of body $b\in\mathcal{B}$ relative to $b_{\text{ref}}$, we start from the retargeted reference motion, represented as keyframes $(\mathbf{q}_{\text{motion}},\mathbf{v}_{\text{motion}})$ of generalized positions and velocities. Using forward kinematics, we obtain for each body $b\in\mathcal{B}$ its pose $T_{b,\text{motion}}$ and twist $\mathcal{V}_{b,\text{motion}}$. Then, we anchor the desired pose for body $b$ as $\hat{T}_{b}=T_{\text{anchor}}\,T^{-1}_{b_{\text{ref}}}\,T_{b,\text{motion}}$, where $T_{b_{\text{ref}}}=(p_{b_{\text{ref}}},R_{b_{\text{ref}}})$ is the current pose of the reference body and $T_{\text{anchor}}=(p_{\text{anchor}},R_{\text{anchor}})$ shifts the motion into the character’s local frame by preserving height, aligning yaw, and translating the $xy$ origin under the robot. Specifically, $p_{\text{anchor}}=[p_{b_{\text{ref}}.x},\ p_{b_{\text{ref}}.y},\ p_{b_{\text{ref}}.z,\text{motion}}]$ and $R_{\text{anchor}}=R_{z}(\mathrm{yaw}(R_{b_{\text{ref}}}R^{\top}_{b_{\text{ref}},\text{motion}}))$. The desired twists remain unchanged, i.e., $\hat{\mathcal{V}}_{b}=\mathcal{V}_{b,\text{motion}}$. Note that for the reference body, $\hat{T}_{b_{\text{ref}}}=T_{b_{\text{ref}},\text{motion}}$ is also unchanged.
Robots often have many closely spaced bodies. Tracking all of them is therefore inefficient and often unnecessary. Instead, we select a subset of target bodies $\mathcal{B}_{\text{target}}\subseteq\mathcal{B}$, forming the motion tracking goal as $g_{\text{tracking}}=(\hat{T}_{b_{\text{ref}}},\hat{T}_{b},\hat{\mathcal{V}}_{b}),\ \forall b\in\mathcal{B}_{\text{target}}$.
@@ III-B Observations
We formulate the policy observation space as a single-timestep vector comprising three components. First, we include the joint positions and velocities from the motion, $\mathbf{c}=[\mathbf{q}_{\text{joint,motion}},\mathbf{v}_{\text{joint,motion}}]$, serving solely as phase information. Note that the policy is not intended to track joint values directly. Second, we include the pose tracking error of the reference body (the anchor), $\xi_{b_{\text{ref}}}\in\mathbb{R}^{9}$, which consists of the three-dimensional position error and the first two columns of the rotation error matrix [37]. Since the reference motion is predefined in the world frame, this term implicitly provides orientation for balancing and global position for correcting drift. Third, we include motion-irrelevant observations, including the robot’s root twist expressed in the root frame $^{b_{\text{root}}}\!\mathcal{V}_{b_{\text{root}}}$, joint positions $\mathbf{q}_{\text{joint}}$ and velocities $\mathbf{v}_{\text{joint}}$, and the previous action $\mathbf{a}_{\text{last}}$.
The complete observation space is then $\mathbf{o}=[\mathbf{c},\xi_{b_{\text{ref}}},{}^{b_{\text{root}}}\!\mathcal{V}_{b_{\text{root}}},\mathbf{q}_{\text{joint}},\mathbf{v}_{\text{joint}},\mathbf{a}_{\text{last}}]$. Note that when position drift compensation is unnecessary or reliable state estimation is unavailable, the linear pose tracking error and the linear root twist can be omitted.
We utilize asymmetric actor-critic to boost training efficiency. Specifically, in addition to the policy observation, for critic inputs, we add in relative pose of each body with respective to the reference body, $T^{-1}_{b_{\text{ref}}}\,T_{b,\text{motion}},\forall b\in\mathcal{B}$, to enable the critic to estimate the tracking error of each body directly in cartesian space.
@@ III-C Joint Impedance and Actions
The policy action is designed as normalized joint position setpoints: $\mathbf{q}_{j,t}=\bar{\mathbf{q}}_{j}+\alpha_{j}\mathbf{a}_{j,t}$, where $\bar{\mathbf{q}}_{j}$ is a constant nominal joint configuration, $\alpha_{j}=0.25\frac{\tau_{j,\text{max}}}{k_{p,j}}$, with $\tau_{j,\text{max}}$ representing the maximum allowable joint torque for joint $j$. This heuristic assumes that contacts generally occur around $\bar{\mathbf{q}}_{j}$ and that the robot hardware design ensures the maximum joint torque is proportional to the expected load. At low gains, these setpoints are not intended as desired position targets; rather, they serve as intermediate variables to generate desired torques, and are intentionally not clipped by joint kinematic limits.
@@ III-D Rewards
Optionally, a global tracking reward for the reference body $b_{\text{ref}}$ can be added, following the same structure as $r_{\text{tracking}}$ but using errors ${\mathbf{e}}_{p,b_{\text{ref}}}$ and ${\mathbf{e}}_{R,b_{\text{ref}}}$.
@@ III-E Termination and Reset
An episode terminates under two conditions, indicating either a fall or tracking failure: (1) when the height or orientation (considering only pitch and roll) errors of the reference body $b_{\text{ref}}$ exceed predefined thresholds, or (2) when any height of the end-effector body $b\in\mathcal{B}_{\text{ee}}$ deviates significantly from the reference trajectory.
@@ V-C Real-world Performance
fightAndSports1_subject1 Full -
@@
By successfully executing a wide spectrum of motions, ranging from static to highly dynamic, from previously studied to firstly demonstrated, and from athletic feats to stylized motions, we demonstrate that our framework is not only robust and accurate, but also scalable, expressive, and broadly applicable. To the best of our knowledge, this is the first unified motion tracking system capable of such range and quality without task-specific tuning.
@@ VIII Acknowledgment
We would like to thank Amazon for using its facility. This work is supported in part by NSF CMMI-1944722, The Robotics and AI Institute, and BAIR Humanoid Intelligence Center, and in part by Stanford Institute for Human-Centered AI, Wu-Tsai Human Performance Alliance, NSF FRR-2153854. We would like to thank Kevin Zakka for the thoughtful discussion, Yuman Gao, and Zhi Su for their help in the experiments.

### v2

@@ Abstract
Learning skills from human motions offers a promising path toward generalizable policies for whole-body humanoid control, yet two key cornerstones are missing: (1) a high-quality motion tracking framework that faithfully transforms large-scale kinematic references into robust and extremely dynamic motions on real hardware, and (2) a distillation approach that can effectively learn these motion primitives and compose them to solve downstream tasks. We address these gaps with BeyondMimic , a real-world framework to learn from human motions for versatile and naturalistic humanoid control via guided diffusion. Our framework provides a motion tracking pipeline capable of challenging skills such as jumping spins, sprinting, and cartwheels with state-of-the-art motion quality. Moving beyond mimicking existing motions and synthesize novel ones, we further introduce a unified diffusion policy that enables zero-shot task-specific control at test time using simple cost functions. Deployed on hardware, BeyondMimic performs diverse tasks at test time, including waypoint navigation, joystick teleoperation, and obstacle avoidance, bridging sim-to-real motion tracking and flexible synthesis of human motion primitives for whole-body control. https://beyondmimic.github.io/.
@@ I Introduction
We introduce BeyondMimic, a real-world framework to address the above two challenges altogether. BeyondMimic first introduces a motion tracking pipeline that is able to perform highly dynamic motions, such as jumping spins, sprinting, and cartwheels, on real humanoid hardware with state-of-the-art motion quality. Beyond simply mimicking these human motions on humanoids, our framework further presents a guided diffusion policy that synthesizes these motions or skills into a single unified policy that supports zero-shot, task-specific control at test time using only simple cost functions, solving downstream tasks flexibly without additional retraining.
@@
  • Guided Diffusion for Humanoids: Application of loss-guided diffusion to real-world humanoid whole-body control, enabling diverse downstream tasks via simple loss functions at test time.
@@ III Scaleable Motion Tracking
We propose a scalable pipeline for training motion-tracking policies. Using a single MDP configuration and a common set of hyperparameters, the pipeline produces high-quality sim-to-real policies given minutes-long references, demonstrating its scalability towards a diverse set of motions. We introduce the details below. Note that since our tracking framework does not have a history, we omit timestep $t$ in the formulation for clarity. We use the subscript m to denote quantities from the reference motion. Unless otherwise stated, all amounts in this section are expressed in the world frame.
@@ III-A Tracking Objective
We start from the retargeted reference motion, represented as keyframes $(\mathbf{q}_{\text{m}},\mathbf{v}_{\text{m}})$ of generalized positions and velocities. Using forward kinematics, we obtain for each body $b\in\mathcal{B}$ its pose $T_{b,\text{m}}$ and twist $\mathcal{V}_{b,\text{m}}$, where $\mathcal{B}$ is the set of all robot bodies. The goal is to reproduce the reference motion on hardware with high fidelity in global coordinates.
Perturbations for robustness during training and the sim-to-real gap often lead to inevitable global drift. To preserve motion style while allowing such drift, the controller should not track absolute body poses.
Therefore, we select an anchor body $b_{\text{anchor}}\in\mathcal{B}$, typically the root or torso, and anchor the motion reference to the desired tracking objective as follows: For the anchor body, use the reference motion directly: $\hat{T}_{b_{\text{anchor}}}=T_{b_{\text{anchor}},\text{m}}$. For non-anchor bodies $b\in\mathcal{B}\setminus\{b_{\text{anchor}}\}$, the desired pose is computed as $\hat{T}_{b}=T_{\Delta}\,T^{-1}_{b_{\text{anchor}},\text{m}}\,T_{b,\text{m}}$, where $T_{b_{\text{anchor}},\text{m}}$ is the pose of the anchor body in reference motion, and $T_{\Delta}=(p_{\Delta},R_{\Delta})$ with $p_{\Delta}=[p_{b_{\text{anchor}}.x},\ p_{b_{\text{anchor}}.y},\ p_{b_{\text{anchor}}.z,\text{m}}]$ and $R_{\Delta}=R_{z}(\operatorname{yaw}(R_{b_{\text{anchor}}}R^{\top}_{b_{\text{anchor}},\text{m}}))$. This hybrid transform shifts the motion into the robot’s local frame by preserving height, aligning yaw, and translating the $xy$ origin under the robot. Finally, the desired twists remain unchanged, i.e., $\hat{\mathcal{V}}_{b}=\mathcal{V}_{b,\text{m}},\ \forall b\in\mathcal{B}$.
Robots often have many closely spaced bodies. Tracking all of them is therefore inefficient and often unnecessary. Instead, we select a subset of target bodies $\mathcal{B}_{\text{target}}\subseteq\mathcal{B}$, forming the motion tracking objective as $g_{\text{tracking}}=(\hat{T}_{b_{\text{anchor}}},\hat{T}_{b},\hat{\mathcal{V}}_{b}),\ \forall b\in\mathcal{B}_{\text{target}}$.
@@ III-B Observations
We formulate the policy observation space as a single-timestep vector comprising three components. (1) Reference phase. We include the joint positions and velocities from the reference motion, $\mathbf{c}=[\mathbf{q}_{\text{joint,m}},\mathbf{v}_{\text{joint,m}}]$, serving solely as phase information; the policy is not intended to track these joint values directly. (2) Anchor pose-tracking error. We include the pose tracking error of the anchor body, $\xi_{b_{\text{anchor}}}\in\mathbb{R}^{9}$, which consists of the three-dimensional position error and the first two columns of the rotation error matrix [37]. Since the reference motion is predefined in the world frame, this term implicitly provides orientation for balancing and global position for correcting drift. (3) Other Proprioceptions. We include the robot’s root twist expressed in the root frame $^{b_{\text{root}}}\!\mathcal{V}_{b_{\text{root}}}$, joint positions $\mathbf{q}_{\text{joint}}$ and velocities $\mathbf{v}_{\text{joint}}$, and the previous action $\mathbf{a}_{\text{last}}$.
The complete observation space is then $\mathbf{o}=[\mathbf{c},\xi_{b_{\text{anchor}}},{}^{b_{\text{root}}}\!\mathcal{V}_{b_{\text{root}}},\mathbf{q}_{\text{joint}},\mathbf{v}_{\text{joint}},\mathbf{a}_{\text{last}}]$. Note that when position drift compensation is unnecessary or reliable state estimation is unavailable, the linear components may be omitted (i.e., the translational part of $\xi_{b_{\text{anchor}}}$ and the linear component of $^{b_{\text{root}}}\!\mathcal{V}_{b_{\text{root}}}$).
We use an asymmetric actor–critic to boost training efficiency. In addition to the policy observation, the critic also receives per-body relative poses w.r.t. the anchor, $T^{-1}_{b_{\text{anchor}}}\,T_{b},\ \forall b\in\mathcal{B}$, enabling it to estimate body-wise tracking errors directly in Cartesian space.
@@ III-C Joint Impedance and Actions
The policy action is designed as normalized joint position setpoints: $\mathbf{q}_{j,t}=\bar{\mathbf{q}}_{j}+\alpha_{j}\mathbf{a}_{j,t}$, where $\mathbf{a}_{j,t}$ is the policy action output, $\bar{\mathbf{q}}_{j}$ is a constant nominal joint configuration, $\alpha_{j}=0.25\frac{\tau_{j,\text{max}}}{k_{p,j}}$, with $\tau_{j,\text{max}}$ representing the maximum allowable joint torque for joint $j$. This heuristic assumes that contacts generally occur around $\bar{\mathbf{q}}_{j}$ and that the robot hardware design ensures the maximum joint torque is proportional to the expected load. At low gains, these setpoints are not intended as desired position targets; rather, they serve as intermediate variables to generate desired torques, and are intentionally not clipped by joint kinematic limits.
@@ III-D Rewards
Optionally, a global tracking reward for the anchor body $b_{\text{anchor}}$ can be added, following the same structure as $r_{\text{tracking}}$ but using errors ${\mathbf{e}}_{p,b_{\text{anchor}}}$ and ${\mathbf{e}}_{R,b_{\text{anchor}}}$.
@@ III-E Termination and Reset
An episode terminates under two conditions, indicating either a fall or tracking failure: (1) when the height or orientation (considering only pitch and roll) errors of the anchor body $b_{\text{anchor}}$ exceed predefined thresholds, or (2) when any height of the end-effector body $b\in\mathcal{B}_{\text{ee}}$ deviates significantly from the reference trajectory.
@@ V-C Real-world Performance
fightAndSports1_subject4 Full -
fight1_subject2 Full -
fight1_subject3 Full -
fight1_subject5 Full -
@@
By successfully executing a wide spectrum of motions—ranging from static to highly dynamic, from well-studied to newly demonstrated, and from athletic feats to stylized motions—we show that our framework is robust and accurate, and scalable, expressive, and broadly applicable. These results demonstrate a unified motion-tracking system that attains this breadth and quality without task-specific tuning.
@@ VIII Acknowledgment
We would like to thank Amazon for letting us use their facility. This work is supported in part by NSF CMMI-1944722, The Robotics and AI Institute, and BAIR Humanoid Intelligence Center, and in part by Stanford Institute for Human-Centered AI, Wu-Tsai Human Performance Alliance, NSF FRR-2153854. We would like to thank Kevin Zakka for the thoughtful discussion, Yuman Gao, and Zhi Su for their help in the experiments.